# Semantic Energy: Detecting LLM Hallucination Beyond Entropy

### v2

@@ 1 Introduction
Large Language Models (LLMs) have been widely deployed in various aspects of production and daily life, demonstrating strong capabilities in different fields (Schlegel et al., 2025; Xiang et al., 2025). However, LLMs are still prone to being influenced by hallucinations and are prone to generate incorrect answers in situations where they lack knowledge, thus misleading users into making errors (Zhou et al., 2024; Farquhar et al., 2024). Recently, uncertainty estimation has been shown to be a reliable indicator for detecting hallucinations, reflecting the tendency of an LLM to generate hallucinations (Xiao & Wang, 2021; Huang et al., 2024). When the uncertainty of an LLM response is high, it often suggests a greater likelihood that the response is a hallucination, prompting further actions such as self-reflection (Renze & Guven, 2024; Kirchhof et al., 2025), regenerating of answers (Xu et al., 2025), or intervention by human experts (Liu et al., 2025; Hopkins et al., 2025).
Entropy is a commonly used metric for estimating uncertainty in LLM (Cheng et al., 2025; Duan et al., 2024). Similarly to traditional discriminative models, high entropy indicates high uncertainty because it means that the model cannot confidently select a particular outcome. However, due to the nature of natural language, the entropy of a single response cannot accurately reflect the reliability of LLMs. Specifically, even though LLMs may not confidently generate the next token, the semantic meaning of any generated token can still be the same. In such cases, we cannot identify an unreliable response simply attributing to its low probability of being generated. To accurately describe the uncertainty of responses composed of natural language, semantics must be considered.
Semantic entropy (Farquhar et al., 2024) is a typical method to characterize the semantic uncertainty of responses, effectively representing the probability that an LLM generates hallucinations. Given a question, semantic entropy involves sampling multiple responses, clustering them based on their semantic meaning, and then replacing individual responses with clusters to calculate entropy, thus achieving semantic-aware uncertainty characterization. Based on this method, a wide range of downstream applications have been developed, such as guiding Chain-of-Thought (CoT) reasoning (Ye et al., 2025) and parallel thinking (Xu et al., 2025). However, semantic entropy has significant drawbacks stemming from entropy itself: it fails to capture the model’s inherent uncertainty, leading to its ineffectiveness in some scenarios.
@@
Several previous studies have shown that logits exhibit stronger inherent capabilities to characterize uncertainty compared to probabilities, and the magnitude of logits can indicate whether the model has undergone adequate training in a given scenario (Liu et al., 2020; Fu et al., 2025; Zhang et al., 2024). For example, in out-of-distribution (OOD) detection, studies have highlighted that the logit values for in-distribution (InD) samples are significantly higher than those for OOD samples (Liu et al., 2020). Recent work named LogToKU (Ma et al., 2025) points out that probabilities lose the intensity information of logits during normalization, thus limiting their ability to represent the inherent uncertainty of LLM. From this insight, we propose a new method to improve the failure cases of Semantic Entropy, termed Semantic Energy. Specifically, for a given prompt, we first perform multiple response samplings, followed by semantic sampling. When calculating the final uncertainty, rather than relying on probability as in Semantic Entropy, we estimate the response uncertainty based on logits, enabling the estimated uncertainty to reflect the model’s inherent uncertainty. Our proposed metric significantly outperforms Semantic Entropy in evaluating the reliability of LLM responses, particularly in scenarios where Semantic Entropy fails. The main contributions are as follows:
@@ 2.1 Estimating LLM Uncertainty with Token-Level Entropy
Let $\bm{q}$} denote a natural language query provided as input to the LLM. Given the prompt $\bm{q}$}, a single response is generated as:
@@
where $\bm{x}$} is a token sequence of length $T$. The model assigns a conditional probability to each token $x_{t}$ given the previous context and the input query $p(x_{t}\mid x_{<t},{\bm{q}})$, and the token-level entropy for position $t$ is defined as:
@@
where $\mathcal{V}$ is the model vocabulary. A common approach to estimating the uncertainty of the full response is to aggregate the token-level entropy. This can be estimated by computing the average:
@@
where $H_{\mathrm{avg}}({\bm{x}})$ represents the uncertainty of the response, with a high entropy indicating greater uncertainty. Considering that different tokens have varying levels of importance for this response, some works (Duan et al., 2024) consider using a weighted average to increase the weight of critical tokens when calculating the entropy:
$H_{\mathrm{wavg}}({\bm{x}})=\sum_{t=1}^{T}w_{t}H_{t},\quad\text{where }\sum_{t=1}^{T}w_{t}=1.$ (4)
Although token-level entropy provides a fine-grained view of local uncertainty during generation, it may not fully capture the diversity in the meanings of different responses. This motivates a higher-level notion of uncertainty based on the distribution over semantically distinct outputs, leading to the concept of semantic entropy (Kuhn et al., 2024).
@@ 2.2 Semantic Entropy and Response Clustering
The semantic entropy is then given by:
@@ 3.2.2 Specific Implementation in LLMs
For a model parameterized by $\bm{\theta}$ with a vocabulary $\mathcal{V}$, the token-level energy distribution becomes:
$p(x_{t}^{(i)},\bm{\theta})=\frac{e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{Z_{\bm{\theta}}},\quad Z_{\bm{\theta}}=\sum_{x_{t}^{(i)}\in\mathcal{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau},$ (13)
where $Z_{\bm{\theta}}$ indicates the partial function of the model parameterized by $\bm{\theta}$. Under a Bayesian treatment, total uncertainty is obtained by marginalizing over $\bm{\theta}$:
$p(x_{t}^{(i)},\mathcal{D})=\int_{\bm{\theta}}p(x_{t}^{(i)},\bm{\theta})\cdot{\tt p}(\bm{\theta}\mid\mathcal{D})\,d\bm{\theta}.$ (14)
However, we cannot obtain all possible $\bm{\theta}$, but we can only estimate uncertainty under specific models. For any given $\bm{\theta}$, the conditional probability of generating a token from a subset vocabulary $\mathcal{V}\subseteq\mathbb{V}$ is:
$p(Z_{\bm{\theta},t}\mid Z_{t})=\frac{\sum_{x_{t}^{(i)}\in\mathcal{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{\sum_{x_{t}^{(i)}\in\mathbb{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}.$ (15)
Therefore, given the model parameterized by $\bm{\theta}$, Eq. 14 can be approximated as:
$\tilde{p}(x_{t}^{(i)},\mathcal{D})=p(x_{t}^{(i)},\theta)\cdot p(Z_{\bm{\theta},t}\mid Z_{t})=\frac{e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{Z_{t}},$ (16)
where $\tilde{p}(x_{t}^{(i)},\mathcal{D})$ indicates a sampled approximation of $p(x_{t}^{(i)},\mathcal{D})$. Similarly to softmax-based probability modeling, the probability of a complete sequence is the joint probability of all tokens in the entire sequence:
$\tilde{p}({\bm{x}}^{(i)})=\prod_{t=1}^{T_{i}}\tilde{p}(x_{t}^{(i)})=\frac{e^{-\sum_{t=1}^{T_{i}}\tilde{E}_{t}^{(i)}/k\tau}}{\prod_{t=1}^{T_{i}}Z_{t}},\quad\tilde{E}({\bm{x}}^{(i)})=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\tilde{E}_{t}^{(i)},$ (17)
where $\tilde{E}({\bm{x}}^{(i)})$ indicates the average sequence-level energy, which can also be interpreted as the energy per unit volume in thermodynamics, that is, the free energy density (Callen, 1993). To extend to semantic clusters, we treat each cluster ${\mathbb{C}}_{k}$ energy as a scaled joint energy:
$\tilde{E}_{\mathrm{Bolt}}({\mathbb{C}}_{k})=\frac{1}{n}\sum_{{\bm{x}}^{(i)}\in{\mathbb{C}}_{k}}\tilde{E}({\bm{x}}^{(i)}),$ (18)
where $n$ are the sampling times for every question.
For an LLM trained with cross-entropy loss, we represent $E(x_{t}^{(i)},\bm{\theta})$ as the negative value of the logit, that is, $E(x_{t}^{(i)},\bm{\theta})=-z_{\bm{\theta}}(x_{t}^{(i)})$, and $k\tau$ is by default the temperature used during LLM training, that is, $k\tau=1$. The final uncertainty is defined as:
@@
which means that the lower the energy, the lower the uncertainty.
@@ 4.1 Setup
We conduct experiments using the Qwen3‑8B model (Yang et al., 2025), and the ERNIE-21B-A3B model (MOE architecture) (Baidu, 2025). Our primary goal is to highlight the differences between probability-based methods and energy-based approaches. Therefore, we use the semantic entropy (Farquhar et al., 2024) as a baseline.
@@
Experiments are performed on standard open-domain QA datasets in both Chinese and English: the Chinese dataset CSQA (He et al., 2024) and the English dataset TriviaQA (Joshi et al., 2017). To assess whether the estimated uncertainty can capture the risk that the model makes errors, we estimate the AUROC between uncertainty scores and correctness (i.e., whether the answer is correct).
@@ 4.3.1 Results on questions with single cluster
In Table 2, we present the case where all responses share the same semantics, that is, all responses are clustered into a single group as described in Sec. 3.1. In this scenario, semantic entropy completely fails, whereas semantic energy is still able to provide a certain level of distinction, resulting in semantic energy achieving an average performance improvement of more than 13% compared in terms of AUROC to semantic entropy in cases where the latter is confident. It is important to note that the value of semantic entropy in such cases is always zero, meaning that its performance reflects the expected behavior when the uncertainty indicator is meaningless, for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
4.3.2 Results on think Mode
We conduct experiments on Qwen-8B using the CSQA dataset and explore the case where the think mode is enabled. Specifically, we activate the think mode but discard the content within <think>...</think> during evaluation, considering only the response portion. The final results, shown in Fig. 2, are consistent with the observations in Table 1.
(a) Semantic Entropy
(b) Semantic Energy
@@ 4.3.3 Advantages of semantic cluster
Inspired by semantic entropy, we incorporate semantic influence when computing energy. If semantics are not considered and the energy of a single response is directly used to characterize the reliability of an LLM’s reply, such as in LogTokU (Ma et al., 2025), a clear problem arises: a single response having high energy does not necessarily mean that the entire cluster of responses sharing the same semantics also has high energy. This is because different responses belonging to the same semantic cluster can still have varying energy values. Therefore, we represent the energy of a response by the energy of the cluster to which it belongs. As shown in Fig. 3, we conduct an ablation study on whether to include semantics. The experimental results demonstrate that incorporating semantics significantly improves the accuracy of uncertainty estimation.
@@ 5 Related Work
Recently, numerous uncertainty estimation methods for LLMs have been proposed. These include methods that utilize natural language for uncertainty feedback, including heuristically designed and trained approaches (Tao et al., 2025; Xiong et al., 2023; Lin et al., 2023); methods that estimate uncertainty based on model states, including those leveraging prior knowledge or statistical observations of model states (Kostenok et al., 2023; Li et al., 2025; Liu et al., 2024), or observing changes under perturbations (Zhang et al., 2025b; Gao et al., 2024); and methods that take into account the semantics of the response, including consistency-based uncertainty characterizations (Lyu et al., 2025; Bartsch et al., 2023; Xiao et al., 2025) and approaches that integrate semantics with model states (Kuhn et al., 2024; Grewal et al., 2024).
@@
The utilization of uncertainty estimation is widely applied in both the post-train and inference phases of LLMs. For example, minimizing entropy during the reinforcement learning process helps reduce uncertainty (Zhang et al., 2025a; Agarwal et al., 2025) and encourages exploration of critical positions with higher uncertainty (Zheng et al., 2025; Cheng et al., 2025). In the inference phase, uncertainty can guide the model’s reasoning path (Ye et al., 2025; Wang et al., 2022), guiding RAG (Guo et al., 2025; Chen & Varoquaux, 2025), collaboration (Dey et al., 2025; Kruse et al., 2025; Cui et al., 2025), and indicate when the model should stop or skip thinking (Xu et al., 2025; Zhu et al., 2025).

### v3

@@ 1 Introduction
Large Language Models (LLMs) have been widely deployed in various aspects of production and daily life, demonstrating strong capabilities in different fields (schlegel2025large; xiang2025vision). However, LLMs are still prone to being influenced by hallucinations and are prone to generate incorrect answers in situations where they lack knowledge, thus misleading users into making errors (zhou2024larger; farquhar2024detecting). Recently, uncertainty estimation has been shown to be a reliable indicator for detecting hallucinations, reflecting the tendency of an LLM to generate hallucinations (xiao2021hallucination; huang2024survey). When the uncertainty of an LLM response is high, it often suggests a greater likelihood that the response is a hallucination, prompting further actions such as self-reflection (renze2024self; kirchhof2025self), regenerating of answers (xu2025adaptive), or intervention by human experts (liu2025not; hopkins2025chatbot).
Entropy is a commonly used metric for estimating uncertainty in LLM (cheng2025reasoning; duan2024shifting). Similarly to traditional discriminative models, high entropy indicates high uncertainty because it means that the model cannot confidently select a particular outcome. However, due to the nature of natural language, the entropy of a single response cannot accurately reflect the reliability of LLMs. Specifically, even though LLMs may not confidently generate the next token, the semantic meaning of any generated token can still be the same. In such cases, we cannot identify an unreliable response simply attributing to its low probability of being generated. To accurately describe the uncertainty of responses composed of natural language, semantics must be considered.
Semantic entropy (farquhar2024detecting) is a typical method to characterize the semantic uncertainty of responses, effectively representing the probability that an LLM generates hallucinations. Given a question, semantic entropy involves sampling multiple responses, clustering them based on their semantic meaning, and then replacing individual responses with clusters to calculate entropy, thus achieving semantic-aware uncertainty characterization. Based on this method, a wide range of downstream applications have been developed, such as guiding Chain-of-Thought (CoT) reasoning (ye2025uncertainty) and parallel thinking (xu2025adaptive). However, semantic entropy has significant drawbacks stemming from entropy itself: it fails to capture the model’s inherent uncertainty, leading to its ineffectiveness in some scenarios.
@@
Several previous studies have shown that logits exhibit stronger inherent capabilities to characterize uncertainty compared to probabilities, and the magnitude of logits can indicate whether the model has undergone adequate training in a given scenario (liu2020energy; fu2025r2r; zhang2024sled). For example, in out-of-distribution (OOD) detection, studies have highlighted that the logit values for in-distribution (InD) samples are significantly higher than those for OOD samples (liu2020energy). Recent work named LogToKU (ma2025estimating) points out that probabilities lose the intensity information of logits during normalization, thus limiting their ability to represent the inherent uncertainty of LLM. From this insight, we propose a new method to improve the failure cases of Semantic Entropy, termed Semantic Energy. Specifically, for a given prompt, we first perform multiple response samplings, followed by semantic sampling. When calculating the final uncertainty, rather than relying on probability as in Semantic Entropy, we estimate the response uncertainty based on logits, enabling the estimated uncertainty to reflect the model’s inherent uncertainty. Our proposed metric significantly outperforms Semantic Entropy in evaluating the reliability of LLM responses, particularly in scenarios where Semantic Entropy fails. The main contributions are as follows:
@@ 2.1 Estimating LLM Uncertainty with Token-Level Entropy
Let $\bm{q}$} denote a natural language query provided as input to the LLM. Given the prompt $\bm{q}$}, a single response sequence is generated in an auto-regressive manner. This response can be represented as:
@@
where $\bm{x}$} denotes a complete token sequence of variable length $T$. At each decoding step $t$, the model computes a probability distribution over its entire vocabulary $\mathcal{V}$, assigning a conditional probability $p(x_{t}\mid x_{<t},{\bm{q}})$ to each candidate token given the preceding context and the original query. To quantify the local uncertainty inherent in the model’s predictions at each generation step, we compute the token-level entropy for position $t$, defined formally as:
@@
where $\mathcal{V}$ is the model vocabulary. A higher value of $H_{t}$ indicates a more uniform probability distribution and thus greater uncertainty in the model’s choice for the $t$-th token.
To estimate the overall uncertainty associated with the entire generated response $\bm{x}$}, a common and straightforward strategy is to aggregate these local token-level entropy values. The most prevalent aggregation method is to compute the arithmetic mean across all tokens in the sequence:
@@
where $H_{\mathrm{avg}}({\bm{x}})$ serves as a proxy for the total uncertainty of the response, with a higher average entropy suggesting a more uncertain generation process. However, this approach implicitly assumes that each token contributes equally to the overall uncertainty, which may not hold in practice. Recognizing that different tokens can carry varying levels of importance for the meaning and correctness of the final response, some recent studies (duan2024shifting) have proposed a refinement by employing a weighted average. This method aims to amplify the contribution of critical or pivotal tokens (e.g., those conveying key facts or decisive information) to the final uncertainty score:
$H_{\mathrm{wavg}}({\bm{x}})=\sum_{t=1}^{T}w_{t}H_{t},\quad\text{where }\sum_{t=1}^{T}w_{t}=1,$ (4)
where weights $w_{t}$ can be determined based on heuristic rules or learned mechanisms designed to identify semantically important tokens.
Although token-level entropy offers a fine-grained, local perspective on the uncertainty during the auto-regressive generation process, it possesses an inherent limitation: it operates purely on a syntactic or surface level. Quantifies the model’s hesitation in choosing the next token, but does not necessarily reflect uncertainty over the underlying meaning or intent of the full response. This is because vastly different token sequences can express the same semantic content, while highly similar token-level probability distributions might lead to responses with divergent meanings. Consequently, token-level metrics may not fully capture the diversity in the semantic content of different possible responses. This critical shortcoming motivates the need for a more holistic, higher-level notion of uncertainty that operates on the distribution of semantically distinct outputs, leading to the concept of semantic entropy (kuhn2023semantic).
@@ 2.2 Semantic Entropy and Response Clustering
Finally, semantic entropy ($H_{\mathrm{SE}}$) is computed by applying the standard Shannon entropy formula to this distribution over semantic clusters:
@@ 3.2.2 Specific Implementation in LLMs
For a model parameterized by $\bm{\theta}$, we approximate $E(x_{t}^{(i)})$ as the negative value of the logit, that is:
$\tilde{E}(x_{t}^{(i)})=-z_{\bm{\theta}}(x_{t}^{(i)}),$ (13)
where $z_{\bm{\theta}}(x_{t}^{(i)})$ is the logits value of $x_{t}^{(i)}$, and $\tilde{E}(x_{t}^{(i)})$ is the approximation of ${E}(x_{t}^{(i)})$. The final uncertainty is defined as:
@@
which means that the lower the energy, the lower the uncertainty. By enhancing semantic entropy with LogTokU, we can better characterize the intrinsic uncertainty of LLMs when answering questions, significantly improving the performance of semantic entropy.
@@ 4.1 Setup
We conduct experiments using the Qwen3‑8B model (yang2025qwen3), and the ERNIE-21B-A3B model (MOE architecture) (baidu2025ernie). Our primary goal is to highlight the differences between probability-based methods and energy-based approaches. Therefore, we use the semantic entropy (farquhar2024detecting) as a baseline.
@@
Experiments are performed on standard open-domain QA datasets in both Chinese and English: the Chinese dataset CSQA (he2024chinese) and the English dataset TriviaQA (joshi2017triviaqa). To assess whether the estimated uncertainty can capture the risk that the model makes errors, we estimate the AUROC between uncertainty scores and correctness (i.e., whether the answer is correct).
@@ 4.3.1 Results on questions with single cluster
In Table 2, we present the case where all responses share the same semantics, that is, all responses are clustered into a single group as described in Sec. 3.1. In this scenario, semantic entropy completely fails, whereas semantic energy is still able to provide a certain level of distinction, resulting in semantic energy achieving an average performance improvement of more than 13% compared in terms of AUROC to semantic entropy in cases where the latter is confident. It is important to note that the value of semantic entropy in such cases is always zero, meaning that its performance reflects the expected performance when the uncertainty indicator is meaningless, for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
@@ 4.3.2 Advantages of semantic cluster
Inspired by semantic entropy, we incorporate semantic influence when computing energy. If semantics are not considered and the energy of a single response is directly used to characterize the reliability of an LLM’s reply, such as in LogTokU (ma2025estimating), a clear problem arises: a single response having high energy does not necessarily mean that the entire cluster of responses sharing the same semantics also has high energy. This is because different responses belonging to the same semantic cluster can still have varying energy values. Therefore, we represent the energy of a response by the energy of the cluster to which it belongs. As shown in Fig. 2, we conduct an ablation study on whether to include semantics. The experimental results demonstrate that incorporating semantics significantly improves the accuracy of uncertainty estimation.
@@
4.3.3 Results on think Mode
To validate the performance of the proposed method on the thinking model, we conduct experiments on Qwen-8B using the CSQA dataset and explore the case where the think mode is enabled. Specifically, we activate the think mode but discard the content within <think>...<think> during evaluation, considering only the response portion. The final results, shown in Fig. 3, are consistent with the observations in Table 1. This indicates that even when the LLM undergoes a lengthy thinking process during output generation, its final results can still accurately capture the uncertainty of the model’s responses through logits, thereby reflecting the reliability of the answers. Additionally, we observe that both semantic entropy and semantic energy demonstrate significantly improved uncertainty characterization capabilities in the think mode compared to the performance reported in Table 1. This suggests that the context during the deep thinking process may positively contribute to characterizing the distributional uncertainty of the final responses.
(a) Semantic Entropy
(b) Semantic Energy
@@ 5 Related Work
Recently, numerous uncertainty estimation methods for LLMs have been proposed. These include methods that utilize natural language for uncertainty feedback, including heuristically designed and trained approaches (tao2025revisiting; xiong2023can; lin2023generating); methods that estimate uncertainty based on model states, including those leveraging prior knowledge or statistical observations of model states (kostenok2023uncertainty; li2025language; liu2404uncertainty), or observing changes under perturbations (zhang2025token; gao2024spuq); and methods that take into account the semantics of the response, including consistency-based uncertainty characterizations (lyu2025calibrating; bartsch2023self; xiao2025consistency) and approaches that integrate semantics with model states (kuhn2023semantic; grewal2024improving).
@@
The utilization of uncertainty estimation is widely applied in both the post-train and inference phases of LLMs. For example, minimizing entropy during the reinforcement learning process helps reduce uncertainty (zhang2025right; agarwal2025unreasonable) and encourages exploration of critical positions with higher uncertainty (zheng2025first; cheng2025reasoning). In the inference phase, uncertainty can guide the model’s reasoning path (ye2025uncertainty; wang2022self), guiding RAG (guo2025empowering; chen2025query), collaboration (dey2025uncertainty; kruse2025information; cui2025entropy), and indicate when the model should stop or skip thinking (xu2025adaptive; zhu2025uncertainty).
@@ 6 Discussion and Conclusion
Appendix A Details for Comparison Methods
In this paper, we primarily compare our method with the well-known approach, semantic entropy. Since both methods require response sampling and semantic clustering, we use identical data for these parts. That is, the only difference lies in the final uncertainty calculation process—the responses and clusters are generated from the same set of data. For the TriviaQA dataset, due to its large number of entries, we only estimate results for the first 5000 samples. Note that this is a commonly adopted practice. Additionally, the sampling temperature is set to 0.6, as recommended in the official documentation, and the random seed is set to values from 1 to 10 to sample ten distinct responses.
It is important to note that the value of semantic entropy in table 2 is always zero, meaning that its performance reflects the expected result when the uncertainty indicator is meaningless; for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
Appendix B Prompts for Different Experiments
[…]